1/4th of the elephants are in woods and double of the square root of total numbers is at the mountain bed.the remaining 15 are at the river bank.Find the total number of elephants.
Answer by stanbon(75887)

1/4th of the elephants are in woods and double of the square root of total numbers is at the mountain bed.the remaining 15 are at the river bank.Find the total number of elephants.
-------
Let the total number of elephants be "x";
Woods: (1/4)x
Mountain bed: 2sqrt(x)
Other: 15
------------------
Equation:
(1/4)x + 2sqrt(x) + 15 = x
x + 8sqrt(x) + 60 = 4x
3x - 8sqrt(x) - 60 = 0
------
Let sqrt(x) = w
Substitute to get:
3w^2 -8w -60 = 0
Factor:
(3w+10)(w-6) = 0
---
Positive solution:
w = 6
----
Substitute:
sqrt(x) = w = 6
x = 36 elephants
===================
Cheers,
Stan H.